Find the values of x and y that solve the system equation
-4x + 5y = -10
5x - 4y = 8

Answer by RAY100(1637)

Multiply 1st eqn by 4 and 2nd eqn by 5
.
-16x+20y=-40
25x-20y = 40,,,,,add
.
9x =0
x=0,,,,,,,subst in 1st eqn
.
-16(0) +20y = -40
y=-2
.
check,,,(1) -4(0)+5(-2)=-10,,,,,ok
(2) 5(0) -4(-2)=8,,,,,,,,ok